OpenStudy (juanilet2000):

Given the center point and the radius, what is the equation of the circle? Center (-5,12) r=13 1. (x+5)^2 + (y-12)^2=169 2. (x+5)^2 + (y-12)^2=13 3. (x-5)^2 + (y+12)^2 = 169 4. (x-5)^2 + (y-12)^2 = 13

OpenStudy (3mar):

May I help?

OpenStudy (juanilet2000):

Sure!

OpenStudy (3mar):

thanks

OpenStudy (3mar):

Do you know the general form of the circle equation?

OpenStudy (3mar):

Here you are!

OpenStudy (juanilet2000):

(x-h)^2 + (y-k)^2 = r^2 ?

OpenStudy (juanilet2000):

Thanks! how do I plug the numbers into the equation?

OpenStudy (3mar):

x_o is the x-coordinate of the center y_o is the y-coordinate of the center

OpenStudy (3mar):

and raise r to the power 2 when you plug it in the equation, r^2=13^2

OpenStudy (juanilet2000):

Ohh i think I have it, so the answer is 3?

OpenStudy (3mar):

but x_o=-5 not 5 so (x+5)^2 not (x-5)^2

OpenStudy (juanilet2000):

So the numbers get flipped?

OpenStudy (3mar):

if x_o=-ve, then (x+x_o)^2 if x_o=+ve, then (x-x_o)^2 and y the same

OpenStudy (juanilet2000):

Okay, I see so the answer is 1?

OpenStudy (3mar):

Congratulations! ♥

OpenStudy (3mar):

Thank you for learning!!

OpenStudy (3mar):

Any more questions?